The length and width of a rectangle have ratio 2 is to 3.The perimeter is 20cm.Find the length and the width.
Answer by ewatrrr(10682)


Hi,
l/w = 2/3
l = (2/3)w
question states
20cm = w + w + (2/3)w + (2/3)w
solve for width
20cm =10/3 w
6cm = width, length then is 4cm
checking our answer
2*6cm + 2*4cm = 12cm + 8cm = 20cm